Should print FRAMLEGG on the first two pages, light, outlined. Should print FRAMLEGG on bottom of the page after the second page.

\usepackage[light,firsttwo,outline,bottomafter]{draftcopy}

Test8: This is a test page 1

This is a testpage 1 Test8: This is a test page 2 This is a testpage 2 Test8: This is a test page 3 This is a testpage 3 Test8: This is a test page 4 This is a testpage 4